Find i (the rate per period) and n (the number of periods) for the following annuity.

Monthly deposits of $2,700 are made for 17 years into an annuity that pays 8.9% compounded manually.

i = ? (Type an integer or decimal.)
Answer #1

Soluton :

Since the annuity is compounded monthly the rate per period = 8.9/12 = .7416%

Number of periods = No of years * frequency of the payments =17 * 12 = 204